how do you solve -1=(5+x)/6
Answer by jessica43(140)

First you need to get rid of the fraction on the right side. You can do this by multiplying each side by 6 because it will cancel out the denominator:
-1 = (5+x)/6
6(-1) = ((5+x)/6)(6)
-6 = 5+x
Now subtract 5 from each side:
-6 - 5 = 5+x-5
-11 = x
So your final answer is x = -11.